The critical angle for light wave going from a medium in which wave length is 4000 A to a medium in which wave length is 6000 A is......
 

Grade:7

1 Answers

Arun
6 years ago
we know that
(v1/v2)=(w1/w2)=(u2/u1)
v=velocity in medium
w=wavelength
u=refractive index
=       u2/u1=6000/4000=3/2
=       u1/u2=2/3
=       u2sin C=u1sin90    Cis critical angle
=        sinC=u1/u2
=        C=sin^-1(2/3)